OpenStudy (anonymous):

A line contains the point (–4, 6) and has a slope of 2004-06-02-03-00_files/i0050000.jpg. Write the equation of the line using point-slope form

OpenStudy (anonymous):

slope of -2/3

OpenStudy (anonymous):

\[ \frac{y-(6)}{x-(-4)} = -\frac 23 \]

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

um im confused

OpenStudy (p0sitr0n):

i dont think the image upload worked

OpenStudy (anonymous):

the image is the slope ^^^

OpenStudy (anonymous):

okay first how do i do this

OpenStudy (anonymous):

\[ \frac{y_2-y_1}{x_2-x_1} = m\implies \frac{y-(6)}{x-(-4)} = -\frac 23 \]

OpenStudy (anonymous):

I let \((x_2,y_2) = (x,y)\), \((x_1,y_1)=(-4,6)\), and \(m=-2/3\).

OpenStudy (anonymous):

is it y+6=-2/3(x-4)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

or y-6=-2/3(x+4)

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

yes, but you need to simplify

OpenStudy (anonymous):

no im good

OpenStudy (anonymous):

thank yu